# Measuring Social Norms of Large Language Models

### v2

@@ Abstract
We present a new challenge to examine whether large language models understand social norms. In contrast to existing datasets, our dataset requires a fundamental understanding of social norms to solve. Our dataset features the largest set of social norm skills, consisting of $402$ skills and $12,383$ questions covering a wide set of social norms ranging from opinions and arguments to culture and laws. We design our dataset according to the K-12 curriculum. This enables the direct comparison of the social understanding of large language models to humans, more specifically, elementary students. While prior work generates nearly random accuracy on our benchmark, recent large language models such as GPT3.5-Turbo and LLaMA2-Chat are able to improve the performance significantly, only slightly below human performance. We then propose a multi-agent framework based on large language models to improve the models’ ability to understand social norms. This method further improves large language models to be on par with humans. Given the increasing adoption of large language models in real-world applications, our finding is particularly important and presents a unique direction for future improvements.
@@
Ye Yuan1,2,3, Kexin Tang1,2, Jianhao Shen1,2, Ming ${}^{\dagger}$, Chenguang ${}^{\dagger}$ 1School of Computer Science, Peking University 2National Key Laboratory for Multimedia Information Processing, Peking University 3Peking University-Anker Embodied AI Lab 4Washington University in St. Louis {yuanye_pku,jhshen,mzhang_cs}@pku.edu.cn tkx@stu.pku.edu.cn, chenguangwang@wustl.edu
@@ 4.1 Main Results
[…]
@@ 4.2 Results Analysis
[…]
@@ 4.3 Ablation Study
[…]
@@ 4.4 Comparison with Human
[…]
@@ B.2 Detailed Experimental Analysis
Which is an example of a law? (a) You should get a new toothbrush every six months. (b) You may not drive above the speed limit. (c) You must clean your room before dinner. (d) You should eat more fruits and vegetables.
@@
Input: law Output: Page: Law Summary: Law is a set of rules that are created and are enforceable by social or governmental institutions to regulate behavior, with its precise definition a matter of longstanding debate. It has been variously described as a science and as the art of justice. State-enforced laws can be made by a group legislature or by a single legislator, resulting in statutes; by the executive through decrees and regulations; or established by judges through precedent, usually in common law jurisdictions. Private individuals may create legally binding contracts, including arbitration agreements that adopt alternative ways of resolving disputes to standard court litigation. The creation of laws themselves may be influenced by a constitution, written or tacit, and the rights encoded therein. The law shapes politics, economics, history and society in various ways and also serves as a mediator of relations between people. Legal systems vary between jurisdictions, with their differences analysed in comparative law. Useful: true
@@
Input: “”+“” Output: NA Useful: false
@@
Output: To determine which of the given options is an example of a law, we need to understand the definition of a law. According to the information provided by the Wikipedia search, a law is a set of rules created and enforceable by social or governmental institutions to regulate behavior. Let’s analyze each option: (a) You should get a new toothbrush every six months. This statement is a recommendation or advice, but it does not represent a law. It does not involve any enforcement by social or governmental institutions. (b) You may not drive above the speed limit. This statement represents a law. It sets a rule that prohibits driving above the speed limit, and it is enforceable by traffic laws and law enforcement agencies. (c) You must clean your room before dinner. This statement is a rule or expectation, but it does not represent a law. It is more likely a household rule or expectation rather than a law enforced by social or governmental institutions. (d) You should eat more fruits and vegetables. Similar to option (a), this statement is a recommendation or advice, but it does not represent a law. It does not involve any enforcement by social or governmental institutions. Based on the analysis, the example of a law among the given options is (b) You may not drive above the speed limit. Useful: true
@@
SocialAgent Component Inputs and Outputs Problem What month comes right before February? (a) August (b) January Retrieval Agent Input: month comes right before February Output: Page: Before Sunrise Summary: Before Sunrise is a 1995 romantic drama film directed by Richard Linklater and co-written by Linklater and Kim Krizan… Before Sunrise also explores time, death, and self-discovery. Useful: false Programming Agent Input: “”+“” Output: NA Useful: false Reasoning Agent Output: To determine the month that comes right before February, we can use our knowledge of the calendar. Step 1: Recall the order of the months in a calendar year: January, February, March, April, May, June, July, August, September, October, November, December. Step 2: Identify the month that comes immediately before February in the calendar year. Based on the given choices, the correct answer is (b) January. We can disregard the information from the other agents as it does not provide any relevant information about the order of the months. Useful: true Answer Predicted String: (b) January
@@
In most state legislatures, the upper house is called the Senate and the lower house is called the House of Representatives. Is the House of Representatives bigger or smaller than the Senate? (a) The House of Representatives is smaller. (b) The House of Representatives is bigger.
@@
Input: size of House of Representatives compared to Senate Output: Page: House of Representatives of the Philippines Summary: The House of Representatives of the Philippines (Filipino: Kapulungan ng mga Kinatawan ng Pilipinas… The 19th Congress has 253 congressional districts. Party-list representatives are elected… Useful: true
@@
Input: “”+“” Output: NA Useful: false
@@
Output: To determine whether the House of Representatives is bigger or smaller than the Senate, we can use the information provided and analyze it step-by-step. 1. The question states that in most state legislatures, the upper house is called the Senate and the lower house is called the House of Representatives. 2. We can refer to the information provided by the other agent, which states that the House of Representatives is the lower house of Congress… Therefore, the answer is: (a) The House of Representatives is smaller. Useful: true
@@ Appendix D Summary of Skills
Grade/Skill
Random
LLaMA2-70B-Chat
[…]
GPT3.5-Turbo
[…]
pre-k/choose-the-uppercase-letter-that-matches-c-k-o-p-s-u-v-w-x-z
70
100
100
100
100
pre-k/choose-the-lowercase-letter-that-matches-c-k-o-p-s-u-v-w-x-z
70
100
100
100
100
pre-k/choose-the-lowercase-letter-that-matches-f-i-j-l-m-t-y
50
100
100
100
100
pre-k/choose-the-lowercase-letter-that-matches-a-b-d-e-g-h-n-q-r
76
100
100
100
100
pre-k/choose-the-uppercase-letter-that-matches-a-b-d-e-g-h-n-q-r
76
100
100
100
100
pre-k/choose-the-uppercase-letter-that-matches-f-i-j-l-m-t-y
80
100
100
100
100
kindergarten/choose-the-lowercase-letter-that-matches-c-k-o-p-s-u-v-w-x-z
41
100
100
100
100
kindergarten/choose-the-uppercase-letter-that-matches-a-b-d-e-g-h-n-q-r
57
100
100
100
100
kindergarten/choose-the-lowercase-letter-that-matches-a-b-d-e-g-h-n-q-r
64
100
100
100
100
kindergarten/choose-the-lowercase-letter-that-matches-f-i-j-l-m-t-y
60
100
100
100
100
kindergarten/choose-the-uppercase-letter-that-matches-c-k-o-p-s-u-v-w-x-z
76
100
100
100
100
kindergarten/choose-the-uppercase-letter-that-matches-f-i-j-l-m-t-y
73
100
100
100
100
kindergarten/which-word-does-not-rhyme
17
67
56
65
100
kindergarten/complete-the-sentence-with-the-correct-sight-word
52
100
100
100
100
kindergarten/is-it-a-telling-sentence-or-an-asking-sentence
57
100
100
100
100
kindergarten/complete-the-sentence-with-the-correct-short-vowel-word
40
88
72
100
100
kindergarten/complete-the-rhyme
54
72
71
41
87
kindergarten/capitalize-the-first-letter-of-a-sentence
64
74
100
100
100
kindergarten/who-what-when-where-or-why
52
66
74
84
83
kindergarten/what-am-i
54
100
100
83
100
kindergarten/read-questions-with-short-vowel-words
52
100
100
68
100
kindergarten/find-the-action-verb
81
100
100
100
100
kindergarten/find-the-complete-sentence
56
91
100
100
100
grade-1/describe-the-difference-between-related-words
44
95
84
100
100
grade-1/which-word-does-not-rhyme
25
58
25
65
88
grade-1/to-have-use-the-correct-form
53
100
100
100
100
grade-1/complete-the-sentence-with-a-two-syllable-word
52
100
100
100
100
grade-1/use-actions-and-dialogue-to-understand-characters
48
100
100
100
100
grade-1/complete-the-sentence-with-the-correct-short-vowel-word
54
100
100
100
100
grade-1/complete-the-rhyme
72
64
41
73
98
grade-1/who-what-when-where-or-why
66
74
100
93
100
grade-1/what-am-i
40
100
100
78
100
grade-1/complete-the-sentence-with-the-correct-sight-word-sets-1-2-3
49
100
100
100
100
grade-1/complete-the-sentence-with-the-correct-sight-word-sets-4-5-6-7
36
100
100
100
100
grade-2/which-word-does-not-rhyme
27
75
34
67
87
grade-2/complete-the-sentence-with-the-correct-sight-word
34
100
100
100
100
grade-2/to-have-use-the-correct-form
52
100
100
100
100
grade-2/complete-the-sentence-with-a-two-syllable-word
25
100
100
100
100
grade-2/complete-the-sentence-with-a-three-letter-consonant-blend-word
76
100
100
100
100
grade-2/use-subordinating-conjunctions
62
84
79
83
100
grade-2/use-the-correct-homophone
63
100
93
100
100
grade-2/is-it-a-complete-sentence-or-a-fragment
81
72
90
100
100
grade-2/use-time-order-words
63
86
96
100
100
grade-3/distinguish-characters-points-of-view
44
100
100
100
100
grade-3/determine-the-meanings-of-words-with-greek-and-latin-roots
36
100
100
100
100
grade-3/describe-the-difference-between-related-words
69
100
100
100
98
grade-3/to-have-use-the-correct-form
64
100
100
100
100
grade-3/use-subordinating-conjunctions
64
77
80
71
81
grade-3/use-the-correct-homophone
52
94
100
100
100
grade-3/is-it-a-complete-sentence-or-a-fragment
52
78
66
100
100
grade-3/use-time-order-words
43
95
76
88
75
grade-3/is-it-a-complete-sentence-or-a-run-on
66
85
66
100
100
grade-3/choose-the-synonym
59
100
100
100
100
grade-3/choose-topic-sentences-for-expository-paragraphs
63
100
100
100
100
grade-3/is-the-sentence-a-statement-question-command-or-exclamation
56
100
100
100
100
grade-3/capitalizing-titles
35
59
58
100
83
grade-3/does-the-adjective-tell-you-what-kind-or-how-many
73
100
100
100
100
grade-3/determine-the-meanings-of-similes
64
100
100
100
100
grade-3/revise-the-sentence-using-a-stronger-verb
66
100
100
81
100
grade-3/choose-between-personal-and-reflexive-pronouns
45
82
72
96
86
grade-3/use-the-correct-subject-or-verb
62
54
77
79
100
grade-3/is-the-noun-singular-or-plural
67
100
72
100
100
grade-3/choose-between-subject-and-object-personal-pronouns
63
85
91
86
100
grade-3/is-it-a-complete-sentence-a-fragment-or-a-run-on
48
91
66
100
100
grade-3/use-actions-and-dialogue-to-understand-characters
50
100
100
100
100
grade-3/use-the-correct-article-a-or-an
69
56
41
58
100
grade-3/determine-the-themes-of-myths-fables-and-folktales
61
100
100
100
100
grade-3/is-the-sentence-in-the-past-present-or-future-tense
29
100
91
100
100
grade-3/choose-the-antonym
47
100
100
100
100
grade-3/to-be-use-the-correct-form
61
100
100
100
100
grade-3/identify-text-structures
55
98
100
100
100
grade-3/does-the-adverb-tell-you-how-when-or-where
52
100
81
100
100
grade-3/which-word-is-a-noun
61
76
78
100
100
grade-3/identify-the-irregular-past-tense-ii
54
100
100
100
100
grade-3/read-about-famous-places
55
100
100
100
100
grade-3/use-key-details-to-determine-the-main-idea
52
100
100
100
100
grade-3/which-sentence-is-in-the-regular-past-tense
60
100
100
100
100
grade-3/which-sentence-uses-an-antonym
50
100
84
100
100
grade-3/choose-between-adjectives-and-adverbs
43
89
88
100
100
grade-3/use-action-verbs
48
100
100
100
100
grade-3/pronoun-verb-agreement
73
73
54
99
100
grade-3/greetings-and-closings-of-letters
48
76
88
88
81
grade-3/make-predictions-about-a-story
42
100
100
100
100
grade-3/draw-inferences-from-a-text
52
100
100
100
100
grade-3/identify-base-words-prefixes-and-suffixes
64
81
76
100
100
grade-3/which-sentence-has-the-same-meaning
59
79
100
100
100
grade-3/identify-plurals-singular-possessives-and-plural-possessives
37
100
100
100
100
grade-3/identify-the-authors-purpose-passages
28
100
87
75
100
grade-3/choose-the-text-that-matches-the-writers-purpose
47
100
100
100
100
grade-3/choose-topic-sentences-for-narrative-paragraphs
53
100
100
100
100
grade-3/which-definition-matches-the-sentence
51
100
74
100
100
grade-3/use-greek-and-latin-roots-as-clues-to-the-meanings-of-words
63
100
100
100
100
grade-3/distinguish-facts-from-opinions
50
100
100
100
100
grade-3/determine-the-meanings-of-greek-and-latin-roots
27
100
84
100
100
grade-3/use-coordinating-conjunctions
38
93
71
100
100
grade-3/is-the-subject-singular-or-plural
37
100
100
100
100
grade-3/choose-reasons-to-support-an-opinion
59
56
91
91
91
grade-3/determine-the-meaning-of-a-word-with-a-suffix-review
32
94
100
100
100
grade-3/use-academic-vocabulary-in-context
29
100
100
100
100
grade-3/select-the-words-that-dont-belong
35
46
55
66
64
grade-3/use-adjectives-to-compare
32
74
81
100
100
grade-3/create-compound-sentences
69
36
35
73
66
grade-3/use-adverbs-to-compare
52
100
77
92
82
grade-3/formatting-and-capitalizing-titles
51
55
50
53
53
grade-3/read-poetry
48
100
100
100
100
grade-3/punctuating-dialogue
49
84
80
82
91
grade-3/show-character-emotions-and-traits
45
78
91
91
91
grade-3/use-the-correct-article-a-an-or-the
32
61
64
55
86
grade-3/choose-the-best-transition
45
100
100
100
100
grade-3/is-the-sentence-simple-or-compound
46
88
82
100
100
grade-3/determine-the-meanings-of-words-with-prefixes-and-suffixes-review
66
100
97
100
100
grade-4/distinguish-characters-points-of-view
54
100
100
100
100
grade-4/determine-the-meanings-of-words-with-greek-and-latin-roots
32
100
100
100
100

### v3

@@ Abstract
We present a new challenge to examine whether large language models understand social norms. In contrast to existing datasets, our dataset requires a fundamental understanding of social norms to solve. Our dataset features the largest set of social norm skills, consisting of $402$ skills and $12,383$ questions covering a wide set of social norms ranging from opinions and arguments to culture and laws. We design our dataset according to the K-12 curriculum. This enables the direct comparison of the social understanding of large language models to humans, more specifically, elementary students. While prior work generates nearly random accuracy on our benchmark, recent large language models such as GPT3.5-Turbo and LLaMA2-Chat are able to improve the performance significantly, only slightly below human performance. We then propose a multi-agent framework based on large language models to improve the models’ ability to understand social norms. This method further improves large language models to be on par with humans. Given the increasing adoption of large language models in real-world applications, our finding is particularly important and presents a unique direction for future improvements. The proposed method and dataset are available in https://huggingface.co/datasets/socialdataset2024/social.
@@
Ye Yuan1,2,3, Kexin Tang1,2, Jianhao Shen1,2, Ming Zhang1,2,3, Chenguang Wang4 1School of Computer Science, Peking University 2National Key Laboratory for Multimedia Information Processing, Peking University 3Peking University-Anker Embodied AI Lab 4Washington University in St. Louis {yuanye_pku,jhshen,mzhang_cs}@pku.edu.cn tkx@stu.pku.edu.cn, chenguangwang@wustl.edu
@@ 4.1 Main Results
To examine the social norm understanding, we evaluate models zero-shot on our datasets. The results are shown in Table 2. Notably, the recent LLMs such as LLaMA2-70B-Chat and GPT3.5-Turbo improve the performance of previous models by a large margin on average. We also see that GPT3.5-Turbo performs better than LLaMA2-70B-Chat. We observe similar increases in both social studies and language arts. The most improvements (24.4%) are brought by GPT3.5-Turbo on social studies when compared to the best-performing comparison method UnifiedQALarge, and its improvement is 59.7% over random accuracy. The main reason is that both LLaMA2-70B-Chat and GPT3.5-Turbo are enhanced with reinforcement learning with human feedback (RLHF) Ouyang et al. (2022), which is designed to align model responses with human values. This is important to social norm understanding. This is clear from the performance comparison between LLaMA2-70B-Chat and LLaMA2-70B in Figure 6. LLaMA2-70B-Chat improves the performance over LLaMA2-70B by 38.6% on average, and the only notable difference is that LLaMA2-70B-Chat is equipped with the RLHF. This also adds explanations about having more improvements in social studies compared to those in language arts. RLHF mainly brings social perspectives without much emphasis on fundamental language phenomena.
Importantly, SocialAgent consistently performs the best on both subjects (Table 2). With SocialAgent, both LLaMA2-70B-Chat and GPT3.5-Turbo improve their performance on our dataset. The best performance is achieved by SocialAgentGPT3.5-Turbo. This shows that our proposed method is able to integrate important social norm knowledge into LLMs.
@@ 4.2 Results Analysis
We show the accuracy of comparison models and our method on the skill level in Figure 7. We show the results on an uncurated list of skills from both subjects. The complete results on the full skill set are in the appendix. We find similar trends with the overall results on corresponding subjects. Recent advancements in LLMs, in particular, RLHF, improve the performance significantly over the previous near random accuracy. With SocialAgent, both GPT3.5-Turbo and LLaMA2-70B-Chat improve their social norm understanding across different social norm skills. GPT3.5-Turbo outperforms LLaMA2-70B-Chat, and SocialAgentGPT3.5-Turbo obtains the best overall performance. While there are social norm skills such as “purpose-of-government” and “reasons-for-opinion” have been mastered by LLMs, there are still plenty of skills such as “the-american-history” and “use-guide-words” remain unlearned, presenting room for further improvements.
@@ 4.3 Ablation Study
To investigate the importance of each key component of SocialAgent  we show the ablation results of LLaMA2-70B-Chat with SocialAgent on the social studies subset in Table 3. The three settings present removing the retrieval agent (“w/o Retrieval Agent”), the programming agent (“w/o Programming Agent”), and the reasoning agent (“w/o Reasoning Agent”) from SocialAgent respectively. Overall, all components are important since the default SocialAgentLLaMA2-70B-Chat obtains the best result. The most significant decrease is brought by removing the reasoning agent. This means that the model has learned certain fundamentals of social norms during their training. So, a better way to prompt the model to obtain the most relevant knowledge is necessary. Moreover, both the retrieval agent and programming agent are essential to incorporate important social norm knowledge into the models.
@@ 4.4 Comparison with Human
It is important to compare models’ social norm understanding to that of humans. We compare the exam scores of both GPT3.5-Turbo and LLaMA2-70B-Chat and our methods with millions of elementary student users of the IXL platform. The results are shown in Figure 9. Overall, both models still underperform average elementary students in terms of understanding social norms. SocialAgent helps improve these models to be on par with human performance. For instance, SocialAgentGPT3.5-Turbo outperforms humans by an average of 0.8%. This result is significant although more advancements are needed to compete with human experts.
@@ B.2 Detailed Experimental Analysis
Which is an example of a law?
(a) You should get a new toothbrush every six months.
(b) You may not drive above the speed limit.
(c) You must clean your room before dinner.
(d) You should eat more fruits and vegetables.
@@
Input: law
Output: Page: Law
Summary: Law is a set of rules that are created and are enforceable by social or governmental institutions to regulate behavior, with its precise definition a matter of longstanding debate. It has been variously described as a science and as the art of justice. State-enforced laws can be made by a group legislature or by a single legislator, resulting in statutes; by the executive through decrees and regulations; or established by judges through precedent, usually in common law jurisdictions. Private individuals may create legally binding contracts, including arbitration agreements that adopt alternative ways of resolving disputes to standard court litigation. The creation of laws themselves may be influenced by a constitution, written or tacit, and the rights encoded therein. The law shapes politics, economics, history and society in various ways and also serves as a mediator of relations between people.
Legal systems vary between jurisdictions, with their differences analysed in comparative law.
Useful: true
@@
Input: “”+“”
Output: NA
Useful: false
@@
Output: To determine which of the given options is an example of a law, we need to understand the definition of a law. According to the information provided by the Wikipedia search, a law is a set of rules created and enforceable by social or governmental institutions to regulate behavior.
Let’s analyze each option:
(a) You should get a new toothbrush every six months. This statement is a recommendation or advice, but it does not represent a law. It does not involve any enforcement by social or governmental institutions.
(b) You may not drive above the speed limit. This statement represents a law. It sets a rule that prohibits driving above the speed limit, and it is enforceable by traffic laws and law enforcement agencies.
(c) You must clean your room before dinner. This statement is a rule or expectation, but it does not represent a law. It is more likely a household rule or expectation rather than a law enforced by social or governmental institutions.
(d) You should eat more fruits and vegetables. Similar to option (a), this statement is a recommendation or advice, but it does not represent a law. It does not involve any enforcement by social or governmental institutions.
Based on the analysis, the example of a law among the given options is (b) You may not drive above the speed limit.
Useful: true
@@
What month comes right before February?
(a) August
(b) January
@@
Input: month comes right before February
Output: Page: Before Sunrise
Summary: Before Sunrise is a 1995 romantic drama film directed by Richard Linklater and co-written by Linklater and Kim Krizan… Before Sunrise also explores time, death, and self-discovery.
Useful: false
@@
Input: “”+“”
Output: NA
Useful: false
@@
Output: To determine the month that comes right before February, we can use our knowledge of the calendar.
Step 1: Recall the order of the months in a calendar year: January, February, March, April, May, June, July, August, September, October, November, December.
Step 2: Identify the month that comes immediately before February in the calendar year.
Based on the given choices, the correct answer is (b) January.
We can disregard the information from the other agents as it does not provide any relevant information about the order of the months.
Useful: true
Answer
Predicted String: (b) January
SocialAgent Component Inputs and Outputs
Problem
In most state legislatures, the upper house is called the Senate and the lower house is called the House of Representatives. Is the House of Representatives bigger or smaller than the Senate?
(a) The House of Representatives is smaller.
(b) The House of Representatives is bigger.
Retrieval Agent
Input: size of House of Representatives compared to Senate
Output: Page: House of Representatives of the Philippines
Summary: The House of Representatives of the Philippines (Filipino: Kapulungan ng mga Kinatawan ng Pilipinas… The 19th Congress has 253 congressional districts. Party-list representatives are elected… Useful: true
Programming Agent
Input: “”+“”
Output: NA
Useful: false
Reasoning Agent
Output: To determine whether the House of Representatives is bigger or smaller than the Senate, we can use the information provided and analyze it step-by-step.
1. The question states that in most state legislatures, the upper house is called the Senate and the lower house is called the House of Representatives.
2. We can refer to the information provided by the other agent, which states that the House of Representatives is the lower house of Congress…
Therefore, the answer is: (a) The House of Representatives is smaller.
Useful: true
@@ Appendix D Summary of Skills
Grade/Skill Random LLaMA2-70B-Chat SocialAgentLLaMA2-70B-Chat GPT3.5-Turbo SocialAgentGPT3.5-Turbo
pre-k/choose-the-uppercase-letter-that-matches-c-k-o-p-s-u-v-w-x-z 70 100 100 100 100
pre-k/choose-the-lowercase-letter-that-matches-c-k-o-p-s-u-v-w-x-z 70 100 100 100 100
pre-k/choose-the-lowercase-letter-that-matches-f-i-j-l-m-t-y 50 100 100 100 100
pre-k/choose-the-lowercase-letter-that-matches-a-b-d-e-g-h-n-q-r 76 100 100 100 100
pre-k/choose-the-uppercase-letter-that-matches-a-b-d-e-g-h-n-q-r 76 100 100 100 100
pre-k/choose-the-uppercase-letter-that-matches-f-i-j-l-m-t-y 80 100 100 100 100
kindergarten/choose-the-lowercase-letter-that-matches-c-k-o-p-s-u-v-w-x-z 41 100 100 100 100
kindergarten/choose-the-uppercase-letter-that-matches-a-b-d-e-g-h-n-q-r 57 100 100 100 100
kindergarten/choose-the-lowercase-letter-that-matches-a-b-d-e-g-h-n-q-r 64 100 100 100 100
kindergarten/choose-the-lowercase-letter-that-matches-f-i-j-l-m-t-y 60 100 100 100 100
kindergarten/choose-the-uppercase-letter-that-matches-c-k-o-p-s-u-v-w-x-z 76 100 100 100 100
kindergarten/choose-the-uppercase-letter-that-matches-f-i-j-l-m-t-y 73 100 100 100 100
kindergarten/which-word-does-not-rhyme 17 67 56 65 100
kindergarten/complete-the-sentence-with-the-correct-sight-word 52 100 100 100 100
kindergarten/is-it-a-telling-sentence-or-an-asking-sentence 57 100 100 100 100
kindergarten/complete-the-sentence-with-the-correct-short-vowel-word 40 88 72 100 100
kindergarten/complete-the-rhyme 54 72 71 41 87
kindergarten/capitalize-the-first-letter-of-a-sentence 64 74 100 100 100
kindergarten/who-what-when-where-or-why 52 66 74 84 83
kindergarten/what-am-i 54 100 100 83 100
kindergarten/read-questions-with-short-vowel-words 52 100 100 68 100
kindergarten/find-the-action-verb 81 100 100 100 100
kindergarten/find-the-complete-sentence 56 91 100 100 100
grade-1/describe-the-difference-between-related-words 44 95 84 100 100
grade-1/which-word-does-not-rhyme 25 58 25 65 88
grade-1/to-have-use-the-correct-form 53 100 100 100 100
grade-1/complete-the-sentence-with-a-two-syllable-word 52 100 100 100 100
grade-1/use-actions-and-dialogue-to-understand-characters 48 100 100 100 100
grade-1/complete-the-sentence-with-the-correct-short-vowel-word 54 100 100 100 100
grade-1/complete-the-rhyme 72 64 41 73 98
grade-1/who-what-when-where-or-why 66 74 100 93 100
grade-1/what-am-i 40 100 100 78 100
grade-1/complete-the-sentence-with-the-correct-sight-word-sets-1-2-3 49 100 100 100 100
grade-1/complete-the-sentence-with-the-correct-sight-word-sets-4-5-6-7 36 100 100 100 100
grade-2/which-word-does-not-rhyme 27 75 34 67 87
grade-2/complete-the-sentence-with-the-correct-sight-word 34 100 100 100 100
grade-2/to-have-use-the-correct-form 52 100 100 100 100
grade-2/complete-the-sentence-with-a-two-syllable-word 25 100 100 100 100
grade-2/complete-the-sentence-with-a-three-letter-consonant-blend-word 76 100 100 100 100
grade-2/use-subordinating-conjunctions 62 84 79 83 100
grade-2/use-the-correct-homophone 63 100 93 100 100
grade-2/is-it-a-complete-sentence-or-a-fragment 81 72 90 100 100
grade-2/use-time-order-words 63 86 96 100 100
grade-3/distinguish-characters-points-of-view 44 100 100 100 100
grade-3/determine-the-meanings-of-words-with-greek-and-latin-roots 36 100 100 100 100
grade-3/describe-the-difference-between-related-words 69 100 100 100 98
grade-3/to-have-use-the-correct-form 64 100 100 100 100
grade-3/use-subordinating-conjunctions 64 77 80 71 81
grade-3/use-the-correct-homophone 52 94 100 100 100
grade-3/is-it-a-complete-sentence-or-a-fragment 52 78 66 100 100
grade-3/use-time-order-words 43 95 76 88 75
grade-3/is-it-a-complete-sentence-or-a-run-on 66 85 66 100 100
grade-3/choose-the-synonym 59 100 100 100 100
grade-3/choose-topic-sentences-for-expository-paragraphs 63 100 100 100 100
grade-3/is-the-sentence-a-statement-question-command-or-exclamation 56 100 100 100 100
grade-3/capitalizing-titles 35 59 58 100 83
grade-3/does-the-adjective-tell-you-what-kind-or-how-many 73 100 100 100 100
grade-3/determine-the-meanings-of-similes 64 100 100 100 100
grade-3/revise-the-sentence-using-a-stronger-verb 66 100 100 81 100
grade-3/choose-between-personal-and-reflexive-pronouns 45 82 72 96 86
grade-3/use-the-correct-subject-or-verb 62 54 77 79 100
grade-3/is-the-noun-singular-or-plural 67 100 72 100 100
grade-3/choose-between-subject-and-object-personal-pronouns 63 85 91 86 100
grade-3/is-it-a-complete-sentence-a-fragment-or-a-run-on 48 91 66 100 100
grade-3/use-actions-and-dialogue-to-understand-characters 50 100 100 100 100
grade-3/use-the-correct-article-a-or-an 69 56 41 58 100
grade-3/determine-the-themes-of-myths-fables-and-folktales 61 100 100 100 100
grade-3/is-the-sentence-in-the-past-present-or-future-tense 29 100 91 100 100
grade-3/choose-the-antonym 47 100 100 100 100
grade-3/to-be-use-the-correct-form 61 100 100 100 100
grade-3/identify-text-structures 55 98 100 100 100
grade-3/does-the-adverb-tell-you-how-when-or-where 52 100 81 100 100
grade-3/which-word-is-a-noun 61 76 78 100 100
grade-3/identify-the-irregular-past-tense-ii 54 100 100 100 100
grade-3/read-about-famous-places 55 100 100 100 100
grade-3/use-key-details-to-determine-the-main-idea 52 100 100 100 100
grade-3/which-sentence-is-in-the-regular-past-tense 60 100 100 100 100
grade-3/which-sentence-uses-an-antonym 50 100 84 100 100
grade-3/choose-between-adjectives-and-adverbs 43 89 88 100 100
grade-3/use-action-verbs 48 100 100 100 100
grade-3/pronoun-verb-agreement 73 73 54 99 100
grade-3/greetings-and-closings-of-letters 48 76 88 88 81
grade-3/make-predictions-about-a-story 42 100 100 100 100
grade-3/draw-inferences-from-a-text 52 100 100 100 100
grade-3/identify-base-words-prefixes-and-suffixes 64 81 76 100 100
grade-3/which-sentence-has-the-same-meaning 59 79 100 100 100
grade-3/identify-plurals-singular-possessives-and-plural-possessives 37 100 100 100 100
grade-3/identify-the-authors-purpose-passages 28 100 87 75 100
grade-3/choose-the-text-that-matches-the-writers-purpose 47 100 100 100 100
grade-3/choose-topic-sentences-for-narrative-paragraphs 53 100 100 100 100
grade-3/which-definition-matches-the-sentence 51 100 74 100 100
grade-3/use-greek-and-latin-roots-as-clues-to-the-meanings-of-words 63 100 100 100 100
grade-3/distinguish-facts-from-opinions 50 100 100 100 100
grade-3/determine-the-meanings-of-greek-and-latin-roots 27 100 84 100 100
grade-3/use-coordinating-conjunctions 38 93 71 100 100
grade-3/is-the-subject-singular-or-plural 37 100 100 100 100
grade-3/choose-reasons-to-support-an-opinion 59 56 91 91 91
grade-3/determine-the-meaning-of-a-word-with-a-suffix-review 32 94 100 100 100
grade-3/use-academic-vocabulary-in-context 29 100 100 100 100
grade-3/select-the-words-that-dont-belong 35 46 55 66 64
grade-3/use-adjectives-to-compare 32 74 81 100 100
grade-3/create-compound-sentences 69 36 35 73 66
grade-3/use-adverbs-to-compare 52 100 77 92 82
grade-3/formatting-and-capitalizing-titles 51 55 50 53 53
grade-3/read-poetry 48 100 100 100 100
grade-3/punctuating-dialogue 49 84 80 82 91
grade-3/show-character-emotions-and-traits 45 78 91 91 91
grade-3/use-the-correct-article-a-an-or-the 32 61 64 55 86
grade-3/choose-the-best-transition 45 100 100 100 100
grade-3/is-the-sentence-simple-or-compound 46 88 82 100 100
grade-3/determine-the-meanings-of-words-with-prefixes-and-suffixes-review 66 100 97 100 100
grade-4/distinguish-characters-points-of-view 54 100 100 100 100
grade-4/determine-the-meanings-of-words-with-greek-and-latin-roots 32 100 100 100 100